Routine equipment maintenance was preformed including reverting to the original NGA capillary columns used on the unit because of inconsistent results. Parts were ordered for the ignition section of the atomic absorption unit. Support was given to the JANUS group, who reviewed the chemistry lab database applications. Worn areas on the lab bench top surface were renewed, and cabinets were cleaned, painted, and relabeled. Sandy Dillard was introduced to some of the lab equipment and procedures.
Two network hubs were replaced with wideband fiber optics based units during the initial underway time. The physically larger size of the units slowed the replacement. This upgrade caused problems with Appletalk zones using the FastPath interface to the ethernet system. The network maintenance jobs were done "after hours" to minimize interference with the JANUS group users.
Several computer viruses were discovered and cleared, primarily from SEDCO machines. SEDCO replaced eight of the PCs on their network.
The catwalk heaters were wired and installed above the catwalk area in preparation for the upcoming high-latitude cruises. The electrical department then removed the heaters for storage until they are needed.
No curatorial representative sailed on Leg 174B. Eight cores were collected, tagging basement. Interstitial water samples were taken for the chemistry lab and physical properties and cryogenic magnetometer measurements were made. The interstitial water samples were squeezed and analyzed. Three X-ray diffraction (XRD) samples were taken. The samples and cores will remain on board for shipment to the BCR with the Leg 175 collection.
Support in the labs was mostly routine, addressing problems with the pycnometer in the physical properties lab and the atomic absorption unit in the chemistry lab. A second EPC recorder was installed in the underway lab to replace the 12-kHz Raytheon machine. The 12-kHz transceiver and transducer were investigated as the system failed in deep water. Both copiers required maintenance and were cleaned.
An ongoing project to remove obsolete Totco and DCS sensors and wiring was outlined and initiated.
The cryogenic magnetometer was filled and tuned during the New York port call. A series of experiments with wash cores and a hybrid core barrel with a stainless steel section were conducted to identify possible sources of magnetic overprints on sediment cores. Initial results concluded that the core barrel was not the major contributor to the overprint.
The microscope lab was used primarily by JANUS members who used the computer facilities.
No core samples were processed. Electrical power distribution in the area was modified to dedicate a circuit to the dishwasher; the extension cords were removed.
Physical properties measurements on the eight cores included multisensor track runs, thermal conductivity, index properties, and shear strength. The pycnometer process controller caused some trouble but is presently working. Information and suggested spare parts were requested. The Tech Help for the lab was brought up to date.
Four of the staff volunteered for the METS team and mustered for two fire drills. The library safety video VCR was moved to SEDCO's lounge for their ongoing Quality Through Teamwork indoctrination classes. Two Safety/training wood shop videos, missing for several legs, have not been returned.
Whereas preparation for the port call was fairly hectic, administration of the plan went fairly well. Anticipated difficulty with the Unions proved to be unwarranted. As this was a cruise ship and passenger terminal, the layout was not optimal for normal dockside loading and unloading.
Sets of positions were collected on site to compare the data quality of the two types of GPS receivers, the Omnistar GPS and Ashtech GPS/GLONASS. The second EPC flatbed recorder was installed, replacing the 12-kHz Raytheon flat bed recorder. The Pelagos annotation software was tested with both recorders; a few problems remain. A replacement power supply for the 12-kHz transceiver is expected in the coming shipment.
The hydraulics system associated with the hose handlers was disassembled to replace valves. Hoses and associated fittings were also replaced. Seals in two hose puller motors failed at different times; one was replaced. Planned space modifications and furniture installation for the underway lab was deferred as the furniture was not shipped to New York City as anticipated.
A test seismic line was conducted on 7 August to work with the yellow multichannel array. Various preamplifier and gain values were used to match the analog record of the multichannel array with the single channel analog record. Various lead weight combinations and lengths of cable played out were tried on the multichannel array to achieve comparable depths with the single channel array at the same speed. Initial conclusions are that the real time analog record from a six channel (six phones) combined does not compare with a 60 phone single-channel array.
Once the X-ray fluorescence unit stabilized, a long series of standards were made for major and minor element analysis in preparation for upcoming legs. Familiarization and training began for Bob Olivas in the X-ray lab, focusing on the XRD. There was exposure to other X-ray lab routines and the JANUS data application that is being developed. Three XRD analyses were made of samples taken from Hole 1074A.
To 174B Special Projects and Problems
To 174B Table of Contents